# Implicit State Estimation via Video Replanning

### v1

@@ Abstract
Video-based representations have gained prominence in planning and decision-making due to their ability to encode rich spatiotemporal dynamics and geometric relationships. These representations enable flexible and generalizable solutions for complex tasks such as object manipulation and navigation. However, existing video planning frameworks often struggle to adapt to failures at interaction time due to their inability to reason about uncertainties in partially observed environments. To overcome these limitations, we introduce a novel framework that integrates interaction-time data into the planning process. Our approach updates model parameters online and filters out previously failed plans during generation. This enables implicit state estimation, allowing the system to adapt dynamically without explicitly modeling unknown state variables. We evaluate our framework through extensive experiments on a new simulated manipulation benchmark, demonstrating its ability to improve replanning performance and advance the field of video-based decision-making.
@@ 1 Introduction
Learning from videos has gained significant traction in decision-making, as videos capture rich visual and dynamic information while aligning with how humans acquire knowledge. These properties make them a powerful medium for specifying tasks and learning diverse skills across contexts. Recent work has shown the effectiveness of video-based frameworks in enabling robots to learn behaviors such as object manipulation Li et al. (2024) and navigation Zhang et al. (2024), highlighting the value of video as a flexible and expressive representation.
This paper focuses on video as a planning representation. Given a goal and current observation, video planning systems generate imagined task executions and convert them into robot actions. Unlike symbolic or latent representations, videos naturally encode both perceptual and action information and generalize across tasks and environments. Prior works Chang et al. (2020); Du et al. (2024a; b) leverage these properties to train universal agents using video-based predictions.
@@
A common solution introduces explicit belief models trained in simulation to infer hidden parameters Memmel et al. (2024); Qi et al. (2023). Yet these methods require prior knowledge of relevant parameters and may struggle when interaction data fails to disambiguate uncertainties. In contrast, humans can effectively resolve these challenges by drawing on past experiences and employing adaptive trial-and-error strategies when prior knowledge is lacking. In the context of video planning, failed interactions are not merely setbacks — they offer critical information about the environment and system parameters.
Figure 1: Implicit state estimation via video replanning. Previous video planning methods rely on the first frame (in red). We propose to incorporate past failures (in blue) to improve future plans.
@@
To validate the effectiveness of our framework, we introduce the Meta-World System Identification Benchmark, a new dataset designed to evaluate online adaptation in decision-making across diverse robotic tasks with unknown system parameters. Experiments show that our method allows for rapid adaptation without explicit system identification. Compared to baselines, including RL methods and other methods for video planning, our approach significantly reduces replanning failures and generates more accurate video plans in dynamic, uncertain environments.
@@ 2 Related Works
Videos have emerged as a popular medium for planning in the physical world. Yang et al. (2024b). For planning actions, people have explored using video to train dynamics model Yang et al. (2024a); Ajay et al. (2023); Zhou et al. (2024); Qin et al. (2024) or reward model Ma et al. (2023); Chen et al. (2021a); Escontrela et al. (2023). Another popular approach involves directly predicting video sequence Du et al. (2024a); Ko et al. (2024); Black et al. (2024) or derived representations such as 3D-correspondence or point tracks to guide downstream policy Yuan et al. (2024); Wen et al. (2024); Bharadhwaj et al. (2025). However, most of these frameworks do not consider environmental uncertainty, which limits their ability to efficiently adapt and replan, often resulting in suboptimal performance in such environments.
@@
System identification focuses on learning system dynamics and estimating parameters using experimental data, with early work on input selection and the Fisher Information Matrix Schön et al. (2011); Ljung (1998); Menda et al. (2020). This research laid the groundwork for active identification strategies, which aim to maximize the amount of information gained from system inputs, a key element in classical system identification Hjalmarsson et al. (1996); Lindqvist & Hjalmarsson (2001); Gerencser & Hjalmarsson (2005). Recent advances in the field have applied these methods to real-world systems, such as identifying physical parameters Xu et al. (2019); Kumar et al. (2019); Mavrakis et al. (2020); Gao et al. (2020; 2022); Memmel et al. (2024). Concurrently, Li et al. (2025) has explored verifier-based frameworks that reject infeasible action proposals based on historical input, enabling online adaptation. Our work can be seen as bridging classical system identification and modern video planning method by injecting the ability to implicitly infer physical parameters from past video data to the model.
Imitation learning.
Imitation learning Zare et al. (2024) is a research field of robot learning, aiming to train an agent to behave like an expert, given the dataset of the interaction of the expert and the environment. Imitation learning has been proven to be an effective and data-efficient approach for robot learning Ho & Ermon (2016); Wang et al. (2017); Lee et al. (2021); Lai et al. (2024); Huang et al. (2024). Behavioral cloning (BC; Torabi et al., 2018) formulates imitating an expert as a supervised learning problem and achieves success on various tasks Florence et al. (2021); Chen et al. (2024); Shridhar et al. (2022); Chi et al. (2024); Brohan et al. (2023); Yeh et al. (2025). Imitation learning uses successful expert demonstration, while we aim to leverage failed interactions with environments.
Offline reinforcement learning.
Offline RL assumes datasets composed of state-action-reward sequences Levine et al. (2020); Figueiredo Prudencio et al. (2024). In contrast to online RL, during training, the agent tries to learn to perform the task by leveraging the dataset, without any interaction to the environment. The paradigm of offline RL has existed for dacades. However, with the rise of deep learning, most modern offline RL algorithms utilize deep neural networks to express the value functions and policies Kidambi et al. (2020); Yu et al. (2020); Kumar et al. (2020b); Yu et al. (2021); Yang et al. (2022). In addition, with the development of large language model and transformer, there are also some research about using transformer for reinforcement learning Chen et al. (2021b); Furuta et al. (2022). Unlike offline RL, our framework do not assume dense reward signals.
@@ 3 Problem Formulation
We consider a setting where certain system parameters $\theta$, such as mass, friction, or utility — essential for solving the task — are unknown. Unlike existing work for explicit state estimation, we do not assume access to the parameterization of $\theta$. We only assume that the underlying $\theta$ can be inferred from interaction data. For example, in a bar-pushing task, the center of mass can be implicitly inferred from an interaction video, regardless of whether the push is successful.
We aim to design a system capable of rapidly solving tasks by planning efficiently and leveraging interactions in the environment. For each environment, we assume access to an experience dataset $\mathcal{D}$, consisting of data tuples $d_{i}=(v_{i},o_{i},s_{i})\in\mathcal{D}$, where $v_{i}$ represents an interaction video, $o_{i}$ denotes the object ID, and $s_{i}$ is a binary success indicator. The dataset contains both successful interactions where $s_{i}=1$ and failed interactions where $s_{i}=0$. Our work focuses on evaluating whether the selected plans can guide successful action sequences. We complement our focus on plan evaluation by using a fixed action prediction module to convert videos into actions for execution.
@@ 4 Approach
Our world is full of uncertainty. For example, when trying to move a heavy piece of furniture, you often don’t know its exact weight, how slippery the floor might be, or whether its legs will flex under pressure. As humans, we actively make new plans and use interaction trials to adapt: infer hidden parameters and narrow the space of possible dynamics.
Our framework follows this intuition. In the video planning framework, we propose to leverage prior interaction videos to refine the agent’s belief about the environment through trial-and-error and video-based planning. Shown in Figure 2, given the current scene and past interactions, it retrieves and updates a latent state embedding that encodes a hypothesis of the hidden parameters, then generates a video plan conditioned on this embedding.
Concretely, our video replanning framework maintains two test-time buffers: one for failed interactions and another for failed plans. At each round, it estimates hidden parameters from failed interactions, then generates new candidate plans guided by this estimation (Section 4.1). A rejection strategy (Section 4.2) selects the plan most different from past failures. The chosen plan is then converted into executable actions via an Action Module (Section 4.3), which can be instantiated with inverse dynamics models, goal-conditioned policies, or trajectory-tracking controllers. In our implementation, we adopt the training-free point-tracking method from AVDC Ko et al. (2024). For extensive details of each component, please refer to Appendix C.
Figure 2: Framework overview. (a) Generate and select plans The Video Plan Generator first generates a set of candidate video plans conditioned on first-frame $f_{0}$ and state embedding $e$. The Rejection Module then selects the plan that is least similar to the plans in the Failed Plan Buffer $B_{P}$. (b) Execute plan and update buffers. The Action Module interacts with the environment by following the selected plan $v^{p*}$. If the interaction trial is not successful, we update the Failed Plan Buffer $B_{P}$ and the Failed Interaction Buffer $B_{I}$. (c) Update belief. The state embedding $e$ is updated using the Retrieval Module based on the Failed Interaction Buffer $B_{I}$. Then, we generate the next batch of video plans according to the updated belief.
@@ 4.1 Video-based Planning with State Embeddings
To plan effectively in environments with unknown dynamics, our system constructs a latent representation of hidden parameters, such as mass or friction, through state embeddings. These embeddings act as hypotheses about the environment configuration and serve as conditioning signals for generating scenario-specific rollouts. By combining prior retrieval with test-time refinement, our framework maintains state embeddings as a flexible and adaptive representation of the environment.
@@
Interaction videos from prior executions with the same object are encoded using a context encoder $E$, producing video embeddings $e^{v}_{j}=E(v_{j})$. To ensure consistency across multiple trials of the same object, we preprocess the dataset $\mathcal{D}$ by grouping entries by object ID and selecting a successful trial as the canonical embedding $e^{o}_{j}$. The choice of encoder $E$ defines the latent feature space; we experiment with pretrained vision encoders such as CLIP (Radford et al., 2021) and DINOv2 (Oquab et al., 2024), which yield semantically structured spaces where different environment modes can be distinguished. Please refer to Section D.1 for visualization of such latent space and Section D.2 for details of dataset $\mathcal{D}$ used in our experiments.
Figure 3: Video Plan Generator and Identification Module. We implement the Video Plan Generator and the Identification Module as two video generative models with shared parameters, switched with a binary trigger is_planning. The Video Plan Generator only learns from successful videos in the experience dataset $D$, while the Identification Module is trained with all videos in $D$, including the failures, enabling the state-embedding refining process.
@@
At inference time, the planner requires a state embedding aligned with the current interaction. To obtain such an embedding, we retrieve candidates from prior interactions by comparing the features of the current video against stored embeddings. Instead of deterministically selecting the closest match, we treat the distances as similarity scores, normalize them with a softmax function, and sample a state embedding according to the resulting probability distribution. This stochastic retrieval helps the planner remain robust when multiple past interactions are plausible. However, retrieval alone can still misalign with the ongoing interaction. To address this, we introduce an optimization-based refinement step illustrated in Figure 3: a generative identification module, which shares parameters with the Video Plan Generator, is trained to produce both successful and unsuccessful outcomes. At inference time, we freeze the module’s weights and optimize the embedding to minimize the discrepancy between its generated rollouts and the observed interaction using a standard denoising diffusion loss. This refinement progressively adjusts the retrieved embedding toward the true environment dynamics, providing a more faithful basis for planning.
@@
Once a state embedding is obtained, the Video Plan Generator uses it to predict a sequence of plausible futures conditioned on the current observation. We adopt a diffusion-based planner following AVDC Ko et al. (2024), but replace language-based conditioning with state embeddings. Specifically, the embedding is projected into the token space of CLIP-Text and appended as an additional conditioning token to the 3D U-Net backbone. The generator is trained end-to-end with a denoising diffusion objective, which encourages realistic rollouts under the hypothesized environment. In practice, we generate M=7M{=}7 future frames at $128{\times}128$ resolution and concatenate them with the initial observation to form an 8-frame video plan. These plans capture not only the high-level task intent but also environment-specific details encoded in the embedding, making them a suitable substrate for downstream rejection and action execution.

### v2

@@ Abstract
Video-based representations have gained prominence in planning and decision-making due to their ability to encode rich spatiotemporal dynamics and geometric relationships. These representations enable flexible and generalizable solutions for complex tasks such as object manipulation and navigation. However, existing video planning frameworks often struggle to adapt to failures at interaction time due to their inability to reason about uncertainties in partially observed environments. To overcome these limitations, we introduce a novel framework that integrates interaction-time data into the planning process. Our approach updates model parameters online and filters out previously failed plans during generation. This enables implicit state estimation, allowing the system to adapt dynamically without explicitly modeling unknown state variables. We evaluate our framework through extensive experiments on a novel simulated manipulation task set, demonstrating its ability to improve replanning performance and advance the field of video-based decision-making.
Machine Learning, ICML
@@ 1 Introduction
Learning from videos has gained significant traction in decision-making, as videos capture rich visual and dynamic information while aligning with how humans acquire knowledge. These properties make them a powerful medium for specifying tasks and learning diverse skills across contexts. Recent work has shown the effectiveness of video-based frameworks in enabling robots to learn behaviors such as object manipulation (Li et al., 2024) and navigation (Zhang et al., 2024), highlighting the value of video as a flexible and expressive representation.
This paper focuses on video as a planning representation. Given a goal and current observation, video planning systems generate imagined task executions and convert them into robot actions. Unlike symbolic or latent representations, videos naturally encode both perceptual and action information and generalize across tasks and environments. Prior works (Chang et al., 2020; Du et al., 2024a, b) leverage these properties to train universal agents using video-based predictions.
@@
Figure 1: Implicit state estimation via video replanning. Existing video planning methods rely solely on the first frame (shown in red) and do not utilize previous failure interaction. We propose a framework that can incorporate past failures to inform future plans, as shown in blue. By leveraging test-time interaction feedback, our framework can uncover hidden system dynamics and efficiently solve tasks where the current observation alone is insufficient.
A common solution introduces explicit belief models trained in simulation to infer hidden parameters (Memmel et al., 2024; Qi et al., 2023). Yet these methods require prior knowledge of relevant parameters and may struggle when interaction data fails to disambiguate uncertainties. In contrast, humans can effectively resolve these challenges by drawing on past experiences and employing adaptive trial-and-error strategies when prior knowledge is lacking. In the context of video planning, failed interactions are not merely setbacks — they offer critical information about the environment and system parameters.
@@
To validate the effectiveness of our framework, we evaluate our approach on the Meta-World System Identification Suite, a novel task set for the Meta-World (Yu et al., 2019) environment designed to evaluate online adaptation in decision-making across diverse robotic tasks with unknown system parameters. Experiments show that our method allows for rapid adaptation without explicit system identification. Compared to baselines, including reinforcement learning methods and other methods for video planning, our approach significantly reduces replanning failures and generates more accurate video plans in dynamic, uncertain environments.
@@ 2 Related Works
Videos have emerged as a popular medium for planning in the physical world (Yang et al., 2024b). In this work, we use the term “planning” to refer to synthesizing a continuous, goal-conditioned trajectory in visual space that can be directly executed or used to guide policy learning. This is distinct from symbolic action sequence search in classical AI planning or explicit policy search in POMDP solvers. Existing approaches have explored using video to train dynamics models (Yang et al., 2024a; Ajay et al., 2023; Zhou et al., 2024; Qin et al., 2024) or reward models (Ma et al., 2023; Chen et al., 2021a; Escontrela et al., 2023). Another popular approach involves directly predicting video sequence (Du et al., 2024a; Ko et al., 2024; Black et al., 2024) or derived representations such as 3D-correspondence or point tracks to guide a downstream policy (Yuan et al., 2024; Wen et al., 2024; Bharadhwaj et al., 2025). However, most of these frameworks do not consider environmental uncertainty, which limits their ability to efficiently adapt and replan, often resulting in suboptimal performance in such environments.
@@
System identification focuses on learning system dynamics and estimating parameters using experimental data, with early work on input selection and the Fisher information matrix (Schön et al., 2011; Ljung, 1998; Menda et al., 2020). This research laid the groundwork for active identification strategies, which aim to maximize the amount of information gained from system inputs, a key element in classical system identification (Hjalmarsson et al., 1996; Lindqvist & Hjalmarsson, 2001; Gerencser & Hjalmarsson, 2005). Recent advances in the field have applied these methods to real-world systems, such as identifying physical parameters (Xu et al., 2019; Kumar et al., 2019; Mavrakis et al., 2020; Gao et al., 2020, 2022; Memmel et al., 2024). Concurrently, Li et al. (2025) has explored verifier-based frameworks that reject infeasible action proposals based on historical input, enabling online adaptation. Our work can be seen as bridging classical system identification and modern video planning method by injecting the ability to implicitly infer physical parameters from past video data to the model.
Few-shot adaptation and latent context inference.
A widely adopted abstraction for few-shot adaptation models task variation through latent context or hidden system parameters. Hidden-Parameter MDPs (Doshi-Velez & Konidaris, 2016; Killian et al., 2017) formalize environment variation using unobserved parameters that must be inferred from interaction data. This perspective has been extended in meta-reinforcement learning through latent context representations that are inferred from experience and used to condition policies (Rakelly et al., 2019; Sämundsson et al., 2018). More recent work has pursued scalable and robust latent-context inference, explicitly accounting for uncertainty in few-shot settings (Zhou et al., 2025; Haldar et al., 2023). In robotics, several methods study adaptation under dynamics shifts and embodiment variation, as well as adaptation to changes in physical properties (Kumar et al., 2021; Cho et al., 2024; Lee et al., 2023). In parallel, retrieval-based approaches reuse past interactions to guide specialization through sub-trajectory reuse (Memmel et al., 2025). Our approach also treats system variation as latent, but instead of maintaining beliefs for tuning policies, we optimize our model’s internal representation directly for test-time planning.
Learning from offline data.
Imitation learning and offline reinforcement learning study how to leverage pre-collected datasets, including expert demonstrations and logged experience, to learn control policies (Zare et al., 2024; Ho & Ermon, 2016; Florence et al., 2021; Levine et al., 2020; Figueiredo Prudencio et al., 2024; Kidambi et al., 2020; Yu et al., 2020). Recent work has further expanded these paradigms using expressive sequence models and diffusion-based policies (Lai et al., 2024; Huang et al., 2024; Yeh et al., 2025; Chen et al., 2021b). While we also learn from offline data, we focus on extracting representations of interaction dynamics from videos, including failed trials, rather than directly optimizing policies from state-action or state-action-reward tuples.
@@ 3 Problem Formulation
We consider a setting where certain system or task parameters $\theta$, such as mass, friction, or utility — essential for solving the task — are unknown. Unlike existing work for explicit state estimation, we do not assume access to the parameterization of $\theta$. We only assume that the underlying $\theta$ can be inferred from interaction data. For example, in a bar-pushing task, the center of mass can be implicitly inferred from an interaction video, regardless of whether the push is successful.
We aim to design a system capable of rapidly solving tasks by planning efficiently and leveraging interactions in the environment. For each environment, we assume access to an experience dataset $\mathcal{D}$, consisting of data tuples $d_{i}=(v_{i},o_{i},s_{i})\in\mathcal{D}$, where $v_{i}$ represents an interaction video, $o_{i}$ denotes the object ID that indexes the underlying physical configuration during data collection, and $s_{i}$ is a binary success indicator. The dataset contains both successful interactions where $s_{i}=1$ and failed interactions where $s_{i}=0$. Our work focuses on evaluating whether the selected plans can guide successful action sequences. We enable this evaluation by using a fixed action prediction module to convert videos into executable actions.
@@ 4 Approach
Our world is full of uncertainty. For example, when trying to move a heavy piece of furniture, you often do not know its exact weight, how slippery the floor might be, or whether its legs will flex under pressure. As humans, we actively make new plans and use interaction trials to adapt: infer hidden parameters and narrow the space of possible dynamics.
Our framework follows this intuition. In the video planning framework, we propose to leverage prior interaction videos to refine the agent’s belief about the environment through trial-and-error and video-based planning. Shown in Figure 2, given the current scene and past interactions, it retrieves and updates a latent embedding that implicitly captures the structure of the hidden parameters, then generates a video plan conditioned on this embedding.
Plans vs. interactions.
In this work, we distinguish between plans and interactions because they originate from fundamentally different sources and serve different roles in our framework, even though both are represented as videos.
A plan is generated by the video planner and represents a hypothetical future trajectory based on current observation. Because the planner is trained only on successful executions, it tends to generate videos that resemble successful behavior. However, when its internal representation of the task or system parameters is inaccurate, it may produce implausible or physically incorrect plans (e.g., generating a video of pushing a door that should be pulled instead). Such inconsistencies typically show during execution, for instance, when the door remains stuck.
In contrast, an interaction is a real execution obtained directly from the environment. Interactions, therefore, reflect the true underlying system, although they may contain suboptimal behaviors. Unlike plans, interactions are never physically incorrect in the sense of violating the actual dynamics of the environment.
Concretely, our video replanning framework maintains two test-time buffers: one for failed interactions and another for failed plans. At each round, it maintains an internal latent embedding that is updated based on observed failures, then generates new candidate plans guided by this latent embedding (Section 4.1). A rejection strategy (Section 4.2) selects the plan most different from past failures. The chosen plan is then converted into executable actions via an Action Module (Section 4.3), which can be instantiated with inverse dynamics models, goal-conditioned policies, or trajectory-tracking controllers. In our implementation, we adopt the training-free point-tracking method from AVDC (Ko et al., 2024). Appendix C shows extensive details of each component.
Figure 2: Framework overview. (a) Generate and select plans The Video Plan Generator first generates a set of candidate video plans conditioned on first-frame $f_{0}$ and state embedding $e$. The Rejection Module then selects the plan that is least similar to the plans in the Failed Plan Buffer $B_{P}$. (b) Execute plan and update buffers. The Action Module interacts with the environment by following the selected plan $v^{p*}$. If the interaction trial is not successful, we update the Failed Plan Buffer $B_{P}$ and the Failed Interaction Buffer $B_{I}$. (c) Update belief. The state embedding $e$ is updated using the Retrieval Module, which executes our retrieval and refinement procedures based on the Failed Interaction Buffer $B_{I}$. Then, we generate the next batch of video plans according to the updated belief.
@@ 4.1 Video-based Planning with State Embeddings
To plan effectively under unknown dynamics, our system learns a latent representation from past observations that implicitly captures hidden parameters, such as mass or friction, through state embeddings. These embeddings act as hypotheses about the environment configuration and serve as conditioning signals for generating scenario-specific rollouts. By combining prior retrieval with test-time refinement, our framework maintains state embeddings as a flexible and adaptive representation of the environment.
@@
A state embedding is a single vector associated with an object instance, summarizing its underlying physical properties (e.g., mass, friction, or functionality) as revealed through interaction dynamics. We construct these per-object state embeddings in two stages: (i) encode every interaction video into a latent feature vector, and (ii) select one representative embedding per object ID.
Stage 1: encoding interaction videos. Let the dataset be $\mathcal{D}=\{(v_{i},\,o_{i})\}_{i=1}^{N}$, where each video $v_{i}$ is associated with an object ID $o_{i}$. We encode every interaction video using a pretrained vision encoder $E$, producing video-level embeddings $e_{i}=E(v_{i})$ for all $i=1,\dots,N$. Interaction videos capture how the object responds to applied actions, so the resulting embeddings inherently reflect information about latent physical properties.
Stage 2: constructing per-object state embeddings. To obtain a single representation for each object ID, we group videos in $\mathcal{D}$ by their object ID. For each object ID $o$, we select one successful trial indexed by $\iota$ and define the corresponding canonical embedding as $e^{o}=e_{\,\iota}$. These canonical embeddings serve as compact descriptors of interaction dynamics and are used by our retrieval and planning modules.
Context encoder $E$.
Given the experience dataset $\mathcal{D}$, the context encoder $E$ determines the feature space that the Retrieval Module and the Video Plan Generator operate on. To implement $E$, we experiment with pre-trained vision encoders known to offer strong semantic representations, such as CLIP (Radford et al., 2021) and DINOv2 (Oquab et al., 2024). To encode videos with these image encoders, we uniformly subsample F=8F{=}8 frames from the input video and encode each frame independently, then concatenate the resulting per-frame features to obtain a single-vector video representation.
@@
At inference time, the planner requires a canonical embedding aligned with the current interaction. To obtain such an embedding, we retrieve candidates from prior interactions by comparing the features of the currently observed interaction video against stored embeddings. Instead of deterministically selecting the closest match, we treat the distances as similarity scores, normalize them with a softmax function, and sample a state embedding according to the resulting probability distribution. This stochastic retrieval helps the planner remain robust when multiple past interactions are plausible. Given an incoming interaction video $v$ with feature $e^{v}$, we compute distances
$d_{i}$ $=\|e^{v}-e^{v_{i}}\|_{2}^{2},$ (1)
and form a retrieval distribution
$p_{i}$ $=\frac{\exp(-d_{i}/\tau)}{\sum_{j}\exp(-d_{j}/\tau)}.$ (2)
However, retrieval alone can still misalign with the ongoing interaction. To address this, we introduce an optimization-based refinement step illustrated in Figure 3: a generative identification module, which shares parameters with the Video Plan Generator, is trained to produce both successful and unsuccessful outcomes. At inference time, we freeze the module’s weights and optimize the embedding to minimize the discrepancy between its generated rollouts and the observed interaction using a standard denoising diffusion loss. Concretely, we sample an initial embedding $e^{(0)}=e^{o_{i}}$ with $i\sim p$, and refine it using
$e^{(k+1)}$ $=e^{(k)}-\eta\nabla_{e}\mathcal{L}_{\text{refine}}(e^{(k)};v),$ (3)
$\mathcal{L}_{\text{refine}}(e;v)$ $=\mathbb{E}_{t,\epsilon}\big[\|\epsilon-\epsilon_{\theta}(x_{t}(v),t,e)\|_{2}^{2}\big],$ (4)
where $\epsilon_{\theta}$ denotes the frozen generative identification model shared with the Video Plan Generator.
This refinement progressively adjusts the retrieved embedding toward the true environment dynamics, providing a more faithful basis for planning. While our full pipeline involves both retrieval and refinement processes, the refinement procedure can operate without initializing from a retrieved state embedding. In this case, we initialize $e^{(0)}$ by sampling from a multivariate Gaussian prior with per-dimension mean and variance estimated from all stored canonical embeddings:
$e^{(0)}\sim\mathcal{N}(\mu,\Sigma).$
We report the performance of this refine-from-scratch setting in the main experiments as a baseline.
Figure 3: Video Plan Generator and Identification Module. We implement the Video Plan Generator and the Identification Module as two video generative models with shared parameters, switched with a binary trigger is_planning. The Video Plan Generator only learns from successful videos in the experience dataset $\mathcal{D}$, while the Identification Module is trained with all videos in $\mathcal{D}$, including the failures, enabling the state-embedding refining process.
@@
Once a state embedding is obtained and converted into a canonical embedding, the Video Plan Generator uses it to predict a sequence of plausible futures conditioned on the current observation. We adopt a diffusion-based planner following AVDC (Ko et al., 2024), but replace language-based conditioning with canonical embeddings. Specifically, the embedding is projected into the token space of CLIP-Text and appended as an additional conditioning token to the 3D U-Net backbone. The generator is trained end-to-end with a denoising diffusion objective, which encourages realistic rollouts under the hypothesized environment. In practice, we generate M=7M{=}7 future frames at $128{\times}128$ resolution and concatenate them with the initial observation to form an 8-frame video plan. These plans are then used to guide the action module throughout the next episode. The plans capture not only the high-level task intent but also environment-specific details encoded in the embedding, making them a suitable substrate for downstream rejection and action execution.